Suppose Tony's Corp. produces a differentiated commodity. The company has substantial control over the price of the product. The company's cost function is estimated to be TC = 200 - 20Q + Q2, where Q is the volume per day. The firm also has the following demand equation P = 60 - 2Q, where P represents the price per unit and Q the volume of daily sales.  

How much profit will the company make or how much loss will the company incur?
Answer #1

Suppose Tonys corp produus a differentiated commodity. Also the company has substantial control ovn the price of the productLow To tul revime will be - TR= 60Q - 2Q2 substitutining Q = 13.33 Waget? and to tal TR= 60X13133) - 2 (13.33) = ( 799.8 - 35